can you please help me to solve this equation.
2/[3(z-120)]=40
please give me step by step answer to get the value of "Z"
Thanks
Nasir

Answer by erica65404(394)

The first step is to distribute the [3(z-120)].
3(z-120) = (3*z)-(3*120) = 3z-360
Now the equation is 2/(3z-360) = 40.
The next step is to multiply each side by (3z-360).
(3z-360)*2/(3z-360) = 40*(3z-360)
In (3z-360)*2/(3z-360) the (3z-360)'s cancel out leaving you with 2=40*(3z-360). Now muliply it out.
40*(3z-360) = (40*3z)-(40*360) = 120z-14,400
You now have 2=120z-14,400. Add 14,400 to each side of the equation.
2+14,400 = 120z-14,400+14,400 The -14,400 and the +14,400 cancel out, leaving you with 120z.
The new equation is 14,402 = 120z. Divide each side by 120.
14,402/120 = 120z/120
Round your answer to the nearest ones place.
z=120